### v1

@@ III.3 Fresnel coefficients
The cosmic-ray simulation framework utilizes a global Cartesian coordinate system in which the positions and momenta of the particles, the positions of the receivers and the electric field components $E_{x}$, $E_{y}$ and $E_{z}$ are defined. To apply the Fresnel coefficients we have to construct the orthogonal vectors $\hat{r}$, $\hat{\theta}$ and $\hat{\phi}$ representing a local spherical coordinate system defined by the incoming direction of the ray, within the global Cartesian coordinate system of the simulation framework. This is shown in Figure 5. The Fresnel coefficients are commonly expressed as scaling factors for the S and P electric field components. The P component is the component in the plane of incidence of the ray, i.e. the plane of the ray, while the S component is the component perpendicular to that plane. Since the electric field is perpendicular to the propagation direction and therefore to the unit vector $\hat{r}$, we know that $E_{\text{S}}=E_{\phi}$ and $E_{\text{P}}=E_{\theta}$.
Figure 5: A visualisation of the relation between a global Cartesian coordinate system and a local spherical coordinate system, defined by the incoming direction of the ray.
From ray tracing we find the propagation direction of the ray and, therefore, can construct the vector $\hat{r}=(r_{x},r_{y},r_{z})$ in the Cartesian basis. Since we know that $\hat{\phi}\cdot\hat{r}=0$ and $\hat{\phi}$ is confined to the $xy$-plane, we find
@@
As the Fresnel coefficients are scaling factors for the S and P components of the electric field, the correction given by Equation 6 can be applied directly to the electric field at the receiver, using $\hat{r}$, $\hat{\phi}$ and $\hat{\theta}$ in the local spherical coordinate system at the receiver. There is no need to calculate the electric field and the unit vectors of the corresponding local spherical coordinate system at the point where the ray meets the air-ice boundary. The effect of propagation from the air-ice boundary to the receiver on the electric field is, following Equation 3, limited to an overall decrease in magnitude by a given factor due to the longer geometrical path length $R$, followed by a rotation in the plane of the ray. The only relevant variable at the air-ice boundary is the incidence angle $\theta_{i}$, i.e. the angle between the refracting or reflecting ray and the normal on the interface, which follows directly from ray tracing.
Both the permeabilities in air and ice can be approximated by the permeability of free space $\mu_{0}$, and the coefficients therefore follow the relations given in [45].
Using Snell’s law they can be expressed as
@@
The variables $n_{1}$ and $n_{2}$ refer to the refractive index values of the first and the second medium at the boundary respectively. The ordering of the two media (i.e. air and ice) can differ depending on whether a ray is being transmitted through the air-ice boundary or reflected on the ice-air boundary. It should be noted that for $\theta_{i}$ greater than the angle of total reflection a non-trivial phase shift occurs, which is not included in this work.
@@ III.4 Focusing factor
$F=\sqrt{\frac{n_{Tx}}{n_{Rx}}\frac{R}{\sin{\theta_{Rec}}}\frac{\Delta\theta_{L}}{\Delta z}},,$ (8)
and is calculated for the ice-to-ice direct and indirect ray tracing. Here $n_{Tx}$ and $n_{Rx}$ are the refractive index values at the transmitter and receiver points respectively, $R$ is the geometrical path length of the ray between the two points, $\theta_{Rec}$ is the angle of reception at the receiver point and $\Delta\theta_{L}$ is the difference in ray launch angles ($\theta_{L}$) for when the receiver is moved away from its ‘original’ depth by an amount given by $\Delta z$. In our work, $\Delta z$ was fixed to 111~{}1cm. Following the approach in [46], we limit the factor to $0.5\leq F\leq 2.0$.
@@
III.5 Transition radiation
As shown in [33], the endpoint formalism naturally includes the emission of transition radiation when the calculations of the emission in the first medium and the emission in the second medium are performed separately, which is how it is implemented in FAERIE. The in-air simulation by CoREAS will generate the sudden death emission, while the continuation of the simulation by the GEANT4 module will lead to the sudden appearance emission. The combination of both gives the transition radiation.
@@ IV Simulation Results
Figure 6: The fluence footprint of the simulated cosmic-ray shower at a depth of 100 m for the in-air emission (left), in-ice emission (middle) and the combined emission (right), using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The simulation was performed using 121 antennas in a star-shaped grid with 8 arms and an antenna spacing of 10 m, indicated by the white dots. For the interpolation of the star grid we used the code described in [48].
@@
Figure 7: The electric field components as a function of time for 4 different antennas positioned along the West axis at a depth of 100 m, using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, the red line indicates the in-ice emission. Note that the range on the y-axis is different for every plot.
@@
Figure 8: The electric field components as a function of time for 4 different antennas positioned along the North axis at a depth of 100 m, using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, the red line indicates the in-ice emission. Note that the range on the y-axis is different for every plot.
@@
In this section we discuss simulation results from FAERIE. We simulated a cosmic-ray cascade initiated by a $E_{p}=10^{17}$ eV proton entering the atmosphere at a zenith angle $\theta=0^{\circ}$ and azimuth angle $\phi=0^{\circ}$. We set the geomagnetic field to $\vec{B}=(16.7525\text{ }\mu\text{T},-52.0874\text{ }\mu\text{T})$, which are the values at a latitude of $89.9588^{\circ}$ South following the IGRF12 model [49]. The first component indicates the horizontal component of the magnetic field and defines the geomagnetic North. The second component indicates the vertical component of the magnetic field, and is upwards for negative values. Thinning was applied during the CORSIKA air shower simulation on electromagnetic particles falling below $10^{-6}E_{p}$, using a weight limit of $w=100$ [50, 51]. For hadrons (except $\pi^{0}$’s) and muons the kinetic energy cut-off of was set to $0.3$ GeV. For electrons and photons (including $\pi^{0}$’s) the kinetic energy cut-off of was set to $0.401$ MeV. We use the QGSJETII-04 high energy hadronic interaction model [52] and the UrQMD low energy hadronic interaction model [53, 54]. The altitude of the ice layer was fixed to that of the South Pole, which is $2.835$ km. For the GEANT4 simulation we used the default production cut-off length for gammas, electrons, positrons and protons, which is $1$ mm. Only charged particles with a kinetic energy above $0.1$ MeV were taken into account for the calculation of the radio emission. The longitudinal particle and energy distributions for the in-air particle cascade are shown in Appendix A (Figure 14), together with the energy deposited in the ice by the in-ice particle cascade (Figure 16). More information about the CPU time and job run time of the simulation can be found in Appendix C.
@@
The in-air emission in Figure 6 (left) clearly shows the typical bean-shape which arises due to the interference between the geomagnetic emission and the Askaryan emission, as described in [55]. The in-ice emission is dominated by Askaryan radiation, and therefore a clear Cherenkov-like ring can be seen. The combination of both gives a unique footprint, showing the bean-like shape from the in-air emission close to the shower axis, enclosed by the Cherenkov-like ring from the in-ice emission.
@@
For the in-air emission the West component of the electric field is the strongest, showing that the geomagnetic emission dominates. The Askaryan emission is radially polarized and is therefore clearly visible for antennas located on the North axis (Figure 8), showing a signal in the North component close to the shower core. This component also shows the expected polarization flip when moving from one side of the shower core to the other. The Askaryan emission can also be seen through its interference with the geomagnetic emission in the antennas on the West axis (Figure 7), causing the total signal in the West component to be stronger on the negative side of the axis compared to that on the positive side. The vertical component of the in-air emission is always small, since most of the in-air emission comes from above the ice surface, resulting in mostly horizontally polarized electric fields.
For the in-ice emission we see a strong West component for antennas on the West axis (Figure 7), and a strong North component for antennas on the North axis (Figure 8), showing the radial polarization of the Askaryan emission. The expected polarization flip when moving from one side of the shower core to the other is also clearly visible. In contrast to the in-air emission, the in-ice emission does have significant vertical components, since the in-ice emission is generated much closer to the antennas.
@@ IV.1 In-ice emission without ray tracing
To show the effect of the exponential density profile on the cosmic-ray in-ice particle cascade development and its corresponding radio emission, we also simulated the cosmic-ray cascade using an ice layer with constant density and constant index of refraction. The ice density was fixed to the value of ${}^{3}$, and the corresponding ice refractive index value was fixed to $n=1.35$, i.e. the value at $z=0$ for the exponential profile given in Equation 2.
Figure 9 shows the fluence along the West axis for the in-ice emission using the constant density and index of refraction profiles, compared to the result obtained using the exponential profiles and ray tracing. The simulation was performed using 201 antennas on each axis using an antenna spacing of 1.5 m. As expected, adding ray tracing to the simulation clearly decreases the radius of the Cherenkov ring, as rays bend towards the shower axis when propagating through the ice.
[…]
@@ IV.2 Frequency dependence
Figure 10: The fluence on the North axis (left) and West axis (right) of the simulated cosmic-ray shower at a depth of 100 m for the combined in-air and in-ice emission when applying an 8th order digital Butterworth band-pass filter, for the frequency bands $30-100$ MHz (upper), $100-300$ MHz (middle) and $300-1000$ MHz (lower). The simulation was performed using 201 antennas each on the North and West axis using an antenna spacing of 1.5 m, with a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$.
To study the frequency dependence of the radio emission we applied an 8th order digital Butterworth band-pass filter to the electric field waveforms for three different frequency bands, and calculated the resulting fluence values for 201 antennas each on the North and West axis with an antenna spacing of 1.5 m, shown in Figure 10.
The two outer fluence peaks correspond to the Cherenkov ring from the in-ice emission, while the inner peak shows the bean-shape from the in-air emission. As expected the relative contribution of the in-ice emission at higher frequencies is larger, since due to the higher density, the in-ice particle cascade is more compressed compared to the in-air cascade. In the limit of a point-like charge, coherence is expected over all frequencies. We also clearly see the width of the Cherenkov cone of the in-ice emission decreasing at higher frequencies, as for higher frequencies the condition for coherence is restricted to smaller length scales and therefore only fulfilled closer to the Cherenkov angle.
@@ IV.3 Primary energy dependence
Figure 11: The fluence footprint of the simulated cosmic-ray shower at a depth of 100 m for the in-air emission (left), in-ice emission (middle) and the combined emission (right), using a primary energy of $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The simulation was performed using 121 antennas in a star-shaped grid with 8 arms and an antenna spacing of 10 m, indicated by the white dots. For the interpolation of the star grid we used the code described in [48].
Figure 12: The electric field components as a function of time for 2 different antennas positioned along the West axis at a depth of 100 m, using a primary energy of $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, the red line indicates the in-ice emission. Note that the range on the y-axis is different for every plot.
Figure 11 shows the radio footprint observed by in-ice antennas at 100100100~{}100m depth when the primary energy is increased by an order of magnitude to $E_{p}=10^{18}$ eV. The random number seeds of the simulation with primary energy $E_{p}=10^{17}$ eV were used, in order to minimize the changes due to stochastic processes. The longitudinal particle and energy distributions for the in-air particle cascade are shown in Appendix A (Figure 15), together with the energy deposited in the ice by the in-ice particle cascade (Figure 16). Comparing Figure 11 to Figure 6, we see that the relative contribution from the in-ice emission is now significantly higher. Increasing the energy of the primary particle results in an air shower particle cascade that develops deeper into the atmosphere, which means a larger fraction of the primary energy is still contained within the particle shower core when it reaches the ice. The fluence of the in-air emission has increased by a factor of $100$, which following Equation 9 indicates that the amplitude of the electric fields is proportional to the primary energy, while that of the in-ice emission has increased by a factor of $\sim 250$. The effect of increasing the primary energy on the energy contained within the particle shower core is illustrated by Figures 16 and 17 in Appendix A. Comparing both figures we see that the energy of the in-ice cascade contained within $1$ m of the shower axis has indeed increased by more than a factor of 10.
Figure 12 shows the electric field components of two antennas positioned along the West axis at a depth of $100$ m using a primary energy of $E_{p}=10^{18}$ eV. Comparing these to the two upper panels of Figure 7 confirms this behaviour. Overall the amplitude of the electric field components of the in-air emission has increased by an order of magnitude, which is less than the increase of the amplitude for the in-ice emission on the Cherenkov ring.

### v4

@@ III.3 Fresnel coefficients
The cosmic-ray simulation framework utilizes a global Cartesian coordinate system in which the positions and momenta of the particles; the positions of the receivers; and the electric field components $E_{x}$, $E_{y}$, and $E_{z}$ are defined. To apply the Fresnel coefficients we have to construct the orthogonal vectors $\hat{r}$, $\hat{\theta}$, and $\hat{\phi}$ representing a local spherical coordinate system defined by the incoming direction of the ray, within the global Cartesian coordinate system of the simulation framework. This is shown in Figure 5. The Fresnel coefficients are commonly expressed as scaling factors for the S and P electric field components. The P component is the component in the plane of incidence of the ray, i.e. the plane of the ray, while the S component is the component perpendicular to that plane. Since the electric field is perpendicular to the propagation direction and therefore to the unit vector $\hat{r}$, we know that $E_{\text{S}}=E_{\phi}$ and $E_{\text{P}}=E_{\theta}$.
Figure 5: A visualization of the relation between a global Cartesian coordinate system and a local spherical coordinate system, defined by the incoming direction of the ray.
From ray tracing we find the propagation direction of the ray and, therefore, can construct the vector $\hat{r}=(r_{x},r_{y},r_{z})$ in the Cartesian basis. Since we know that $\hat{\phi}\cdot\hat{r}=0$ and $\hat{\phi}$ is confined to the $xy$ plane, we find
@@
As the Fresnel coefficients are scaling factors for the S and P components of the electric field, the correction given by Equation 6 can be applied directly to the electric field at the receiver, using $\hat{r}$, $\hat{\phi}$, and $\hat{\theta}$ in the local spherical coordinate system at the receiver. There is no need to calculate the electric field and the unit vectors of the corresponding local spherical coordinate system at the point where the ray meets the air-ice boundary. The effect of propagation from the air-ice boundary to the receiver on the electric field is, following Equation 3, limited to an overall decrease in magnitude by a given factor due to the longer geometrical path length $R$, followed by a rotation in the plane of the ray. The only relevant variable at the air-ice boundary is the incidence angle $\theta_{i}$, i.e., the angle between the refracting or reflecting ray and the normal on the interface, which follows directly from ray tracing.
Both the permeabilities in air and ice can be approximated by the permeability of free space $\mu_{0}$, and the coefficients therefore follow the relations given in [48]. Using Snell’s law they can be expressed as
@@
The variables $n_{1}$ and $n_{2}$ refer to the refractive index values of the first and the second medium at the boundary respectively. The ordering of the two media (i.e., air and ice) can differ depending on whether a ray is being transmitted through the air-ice boundary or reflected on the ice-air boundary. It should be noted that for $\theta_{i}$ greater than the angle of total reflection a nontrivial phase shift occurs, which is not included in this work.
@@ III.4 Focusing factor
$F=\sqrt{\frac{n_{Tx}}{n_{Rx}}\frac{R}{\sin{\theta_{Rec}}}\frac{\Delta\theta_{L}}{\Delta z}},$ (8)
and is calculated for the ice-to-ice direct and indirect ray tracing. Here $n_{Tx}$ and $n_{Rx}$ are the refractive index values at the transmitter and receiver points respectively, $R$ is the geometrical path length of the ray between the two points, $\theta_{Rec}$ is the angle of reception at the receiver point, and $\Delta\theta_{L}$ is the difference in ray launch angles ($\theta_{L}$) for when the receiver is moved away from its “original” depth by an amount given by $\Delta z$. In our work, $\Delta z$ was fixed to 111~{}1cm. Following the approach in [49], we limit the factor to $0.5\leq F\leq 2.0$.
@@ IV Simulation Results
Figure 6: The fluence footprint of the simulated cosmic-ray shower at a depth of 100 m for the in-air emission (left), in-ice emission (middle), and the combined emission (right), using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The simulation was performed using 121 antennas in a star-shaped grid with 8 arms and an antenna spacing of 10 m, indicated by the white dots. For the interpolation of the star grid we used the code described in [51].
@@
Figure 7: The electric field components as a function of time for 4 different antennas positioned along the west axis at a depth of 100 m, using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, and the red line indicates the in-ice emission. Note that the range on the y axis is different for every plot.
@@
Figure 8: The electric field components as a function of time for 4 different antennas positioned along the north axis at a depth of 100 m, using a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, and the red line indicates the in-ice emission. Note that the range on the y axis is different for every plot.
@@
In this section we discuss simulation results from FAERIE. We simulated a cosmic-ray cascade initiated by a $E_{p}=10^{17}$ eV proton entering the atmosphere at a zenith angle $\theta=0^{\circ}$ and azimuth angle $\phi=0^{\circ}$. We set the geomagnetic field to $\vec{B}=(16.7525\text{ }\mu\text{T},-52.0874\text{ }\mu\text{T})$, which are the values at a latitude of $89.9588^{\circ}$ south following the IGRF12 model [52]. The first component indicates the horizontal component of the magnetic field and defines the geomagnetic north. The second component indicates the vertical component of the magnetic field, and is upward for negative values. Thinning was applied during the CORSIKA air shower simulation on electromagnetic particles falling below $10^{-6}E_{p}$, using a weight limit of $w=100$ [53, 54]. For hadrons (except $\pi^{0}$’s) and muons the kinetic energy cutoff was set to $0.3$ GeV. For electrons and photons (including $\pi^{0}$’s) the kinetic energy cutoff was set to $0.401$ MeV. We use the QGSJETII-04 high energy hadronic interaction model [55] and the UrQMD low energy hadronic interaction model [56, 57]. The altitude of the ice layer was fixed to that of the South Pole, which is $2.835$ km and corresponds to a vertical atmospheric depth of $729$ g/cm2. For the GEANT4 simulation we used the default production cut-off length for gammas, electrons, positrons and protons, which is $1$ mm. Only charged particles with a kinetic energy above $0.1$ MeV were taken into account for the calculation of the radio emission. The longitudinal particle and energy distributions for the in-air particle cascade are shown in Appendix A (Figure 14), together with the energy deposited in the ice by the in-ice particle cascade (left-hand side of Figure 17). More information about the CPU time and job run time of the simulation can be found in Appendix C.
@@
The in-air emission in Figure 6 (left) clearly shows the typical bean shape which arises due to the interference between the geomagnetic emission and the Askaryan emission, as described in [58]. The in-ice emission is dominated by Askaryan radiation, and therefore a clear Cherenkov-like ring can be seen. The combination of both gives a unique footprint, showing the beanlike shape from the in-air emission close to the shower axis, enclosed by the Cherenkov-like ring from the in-ice emission.
@@
For the in-air emission the west component of the electric field is the strongest, showing that the geomagnetic emission dominates. The Askaryan emission is radially polarized and is therefore clearly visible for antennas located on the north axis (Figure 8), showing a signal in the north component close to the shower core. This component also shows the expected polarization flip when moving from one side of the shower core to the other. The Askaryan emission can also be seen through its interference with the geomagnetic emission in the antennas on the west axis (Figure 7), causing the total signal in the west component to be stronger on the negative side of the axis compared to that on the positive side. The vertical component of the in-air emission is always small since the in-air emission comes from above the ice surface, corresponding to down-going rays toward the given antenna positions and thus horizontally polarized electric fields. It is noticed that the pulses of the west component and the vertical component for the in-air emission at 20 m from the axis center shown in Figure 7 and Figure 8 are different in shape. A possible explanation would be the difference in pulse shapes due to the geomagnetic emission driving the east-west component, and the charge-excess emission driving the vertical component [60].
For the in-ice emission we see a strong west component for antennas on the west axis (Figure 7), and a strong north component for antennas on the north axis (Figure 8), showing the radial polarization of the Askaryan emission. The expected polarization flip when moving from one side of the shower core to the other is also clearly visible. In contrast to the in-air emission, the in-ice emission does have significant vertical components. Since the given antenna positions are located closer to the emission points the incoming rays are not completely down going, which results in electric fields with both a horizontal and vertical component.
@@ IV.1 In-ice emission without ray tracing
To show the effect of the exponential density profile on the cosmic-ray in-ice particle cascade development and its corresponding radio emission, we also simulated the cosmic-ray cascade using an ice layer with constant density and constant index of refraction. The ice density was fixed to the value of 359359359~{}359kg///m3, and the corresponding ice refractive index value was fixed to $n=1.35$, i.e. the value at $z=0$ for the exponential profile given in Equation 2.
Figure 9 shows the fluence along the west axis for the in-ice emission using the constant density and index of refraction profiles, compared to the result obtained using the exponential profiles and ray tracing. In both cases an 8th order digital Butterworth bandpass filter for a frequency band of $300-1000$ MHz was applied. The simulation was performed using 201 antennas on each axis using an antenna spacing of 1.5 m. As expected, adding ray tracing to the simulation clearly decreases the radius of the Cherenkov ring, as rays bend toward the shower axis when propagating through the ice.
Figure 9: The fluence on the west axis of the simulated cosmic-ray shower at a depth of $100$ m for the in-ice emission only using a constant ice density of $359$ kg/m3 and a constant index of refraction $n=1.35$, compared to the result using ray tracing with an exponential density and index of refraction profile. In both cases an 8th order digital Butterworth bandpass filter for a frequency band of $300-1000$ MHz was applied. The simulation was performed using 201 antennas using an antenna spacing of 1.5 m, with a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$.
@@ IV.2 Frequency dependence
Figure 10: The fluence on the north axis (left) and west axis (right) of the simulated cosmic-ray shower at a depth of 100 m for the combined in-air and in-ice emission when applying an 8th order digital Butterworth bandpass filter, for the frequency bands $30-100$ MHz (upper), $100-300$ MHz (middle) and $300-1000$ MHz (lower). The simulation was performed using 201 antennas each on the north and west axis using an antenna spacing of 1.5 m, with a primary energy of $E_{p}=10^{17}$ eV and zenith angle $\theta=0$.
To study the frequency dependence of the radio emission we applied an 8th order digital Butterworth bandpass filter to the electric field waveforms for three different frequency bands, and calculated the resulting fluence values for 201 antennas each on the north and west axis with an antenna spacing of 1.5 m, shown in Figure 10.
The two outer fluence peaks correspond to the Cherenkov ring from the in-ice emission, while the inner peak shows the bean shape from the in-air emission. As expected the relative contribution of the in-ice emission at higher frequencies is larger, since due to the higher density, the in-ice particle cascade is more compressed compared to the in-air cascade. In the limit of a pointlike charge, coherence is expected over all frequencies. We also clearly see the width of the Cherenkov cone of the in-ice emission decreasing at higher frequencies, as for higher frequencies the condition for coherence is restricted to smaller length scales and therefore only fulfilled closer to the Cherenkov angle [14].
@@ IV.3 Primary energy dependence
Figure 11: The fluence footprint of the simulated cosmic-ray shower at a depth of 100 m for the in-air emission (left), in-ice emission (middle), and the combined emission (right), using a primary energy of $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The simulation was performed using 121 antennas in a star-shaped grid with 8 arms and an antenna spacing of 10 m, indicated by the white dots. For the interpolation of the star grid we used the code described in [51].
Figure 12: The electric field components as a function of time for 2 different antennas positioned along the west axis at a depth of 100 m, using a primary energy of $E_{p}=10^{18}$ eV and zenith angle $\theta=0$. The coordinate of each antenna is indicated at the top of the plots. The blue line indicates the in-air emission, and the red line indicates the in-ice emission. Note that the range on the y axis is different for every plot.
Figure 11 shows the radio footprint observed by in-ice antennas at 100100100~{}100m depth when the primary energy is increased by an order of magnitude to $E_{p}=10^{18}$ eV. The longitudinal particle and energy distributions for the in-air particle cascade are shown in Appendix A (Figure 15), together with the energy deposited in the ice by the in-ice particle cascade (right of Figure 17).
The number of particles in the cascade and therefore the amplitude of the electric field scales with primary energy. Following Equation 9, increasing the energy of the primary particle by a factor of $10$ should thus lead to an increase in the fluence by a factor of $100$. However, comparing Figure 11 to Figure 6, we see that the fluence from the in-ice emission at a primary energy of $10^{18}$ eV has increased significantly more compared to that of the in-air emission. An air shower particle cascade with a higher primary energy will develop deeper into the atmosphere, which means a larger fraction of the primary energy is still contained within the particle shower core when it reaches the ice. Increasing the primary energy of the shower by a factor of $10$ has lead to an increase in the total energy deposited in the ice within a radius of $1$ m of the shower axis by a factor of $15$. The fluence integrated over the footprint in the ice of the in-ice emission has increased by a factor of $240$, which is indeed close to the expected $15^{2}$. As a larger part of the shower development has been shifted into the ice, the in-air emission has only increased by a factor of $85$. The effect of increasing the primary energy on the energy contained within the particle shower core is illustrated by Figure 17 in Appendix A.
Figure 12 shows the electric field components of two antennas positioned along the west axis at a depth of $100$ m using a primary energy of $E_{p}=10^{18}$ eV. Comparing these to the two upper panels of Figure 7 confirms this behavior. Overall the amplitude of the electric field components of the in-air emission has increased by 1 order of magnitude, which is less than the increase of the amplitude for the in-ice emission on the Cherenkov ring.